OpenStudy (anonymous):

(FTOC) Evaluate the integral: (x-1)/x^(1/2) dx, [9,1] Answer: -40/3 Me: 12 (x-1)/x^(1/2) dx = (1/2)x^(3/2) - (1/2)x^(1/2) so: [(1/2)(9)^(3/2) - (1/2)(9)^(1/2)] - [(1/2)(1)^(3/2) - (1/2)(1)^(1/2) = [(27/2)-(3/2)] - [0] =12 I must be doing something wrong on doing the (1/2)x^(3/2) - (1/2)x^(1/2) but I cannot figure it out.

zepdrix (zepdrix):

\[\huge \int\limits_1^9 \frac{x-1}{x^{1/2}}dx= \int\limits_1^9 \frac{x}{x^{1/2}}-\frac{1}{x^{1/2}}dx\]

zepdrix (zepdrix):

\[\huge \int\limits_1^9 x^{1/2}-x^{-1/2}dx\]

zepdrix (zepdrix):

Everything look ok so far? :o

OpenStudy (anonymous):

ok I can follow the first one

OpenStudy (anonymous):

what happened to the x above x^1/2

zepdrix (zepdrix):

Remember when you divide terms that have similar bases, you SUBTRACT the exponents :D This is what that should look like. \[\huge \frac{x^1}{x^{1/2}}=x^{1-\frac{1}{2}}=x^{1/2}\]

OpenStudy (anonymous):

duh, yes now both make sense.

zepdrix (zepdrix):

ah good c:

OpenStudy (anonymous):

\[\frac{ x ^{1/2} + ^ {1} }{ 1 }\]

zepdrix (zepdrix):

The equation tool is a little tricky ^^ takes some time to get used to it

zepdrix (zepdrix):

Woops, for the power rule of Integration, you raise the power, then divide by the NEW power. So your new power on the bottom should be 1/2 + 1

zepdrix (zepdrix):

\[\huge \int\limits x^n dx = \frac{x^{n+1}}{n+1}\]\[\huge \int\limits x^{1/2}dx=\frac{x^{(1/2+1)}}{1/2+1}\]

OpenStudy (anonymous):

so \[\frac{ 3 }{ 2 }(x^{3/2})?\]

zepdrix (zepdrix):

We're DIVIDING by 3/2, remember what we do when we divide by a fraction?

OpenStudy (anonymous):

Grrrrr ....... Heh it's the Algebra that kills ya. 2/3's

zepdrix (zepdrix):

heh c: ya

OpenStudy (anonymous):

\[- 2(x^{1/2})\]

OpenStudy (anonymous):

for the 2nd part.

zepdrix (zepdrix):

Mmmm yah that looks good.

OpenStudy (anonymous):

cool .... now plug 9 into both and subtract plugging 1 into both and that should be my new answer.

zepdrix (zepdrix):

yah sounds good c:

OpenStudy (anonymous):

thanks again.